Project

General

Profile

Task #24877

DR_1_Run_8

Added by Daniil Frolov about 2 months ago. Updated 22 days ago.

Status:
New
Priority:
Normal
Assignee:
-
Start date:
08/27/2020
Due date:
% Done:

0%

Estimated time:
Duration:

Description

Initial information on DR1_Run_8.

Main goal: measure Rigetti qubits inside 3 GHz cavities.

Schematics of the fridge can be found here:
https://cdcvs.fnal.gov/redmine/projects/runs/repository/dr1_run_8/revisions/master/raw/DR1_run_8.pdf

Warm electronics and warm switches schematics is located here https://cdcvs.fnal.gov/redmine/projects/runs/repository/dr1_run_8/revisions/master/raw/DR1_run_8_warm_RF.pdf

Everything can be controlled from Labber including warm switches, but cryo switches are still controlled manually.

We have several devices installed:

1) RI-1 (TERI001) 3 GHz cavity with Rigetti qubit + eccosorb filter. Cavity is located on the side.
2) RI-2 (TERI002) 3 GHz cavity with Rigetti qubit + eccosorb filter. Cavity is located in the center.
3) Pappas' qubit + added 10 dB more in the last attenuator. (see: https://cdcvs.fnal.gov/redmine/issues/24199 )
4) McDermott's qubit is same as last DR1_run_7.
5) Pappas' 5 GHz JPA (see: https://cdcvs.fnal.gov/redmine/issues/22808#note-18 )
6) Pappas' 3 GHz JPA
7) Lincoln Lab TWPA (see: https://cdcvs.fnal.gov/redmine/issues/22808#note-12 )

Rigetti_version2.pdf (118 KB) Rigetti_version2.pdf Taeyoon Kim, 08/27/2020 05:32 PM
RI-2_mainMode.JPG (78.8 KB) RI-2_mainMode.JPG Daniil Frolov, 08/31/2020 03:22 PM
8312020441.JPG (42.4 KB) 8312020441.JPG Daniil Frolov, 08/31/2020 04:44 PM
decay.csv (26.2 KB) decay.csv decay data (10 averages) Daniil Frolov, 08/31/2020 04:44 PM
RI-1.png (70.5 KB) RI-1.png Daniil Frolov, 08/31/2020 04:59 PM
decay_RI-1.csv (26.2 KB) decay_RI-1.csv full horizontal scale = 1s Daniil Frolov, 08/31/2020 05:14 PM
8312020515.JPG (47 KB) 8312020515.JPG Daniil Frolov, 08/31/2020 05:15 PM
912020328.JPG (125 KB) 912020328.JPG Daniil Frolov, 09/01/2020 03:31 PM
9220201002.JPG (126 KB) 9220201002.JPG Daniil Frolov, 09/02/2020 10:02 AM

History

#1 Updated by Taeyoon Kim about 2 months ago

I checked the GDS file that the Rigetti team sent, and I found the dimensions are a bit modified from what I sent. But the parameters are not hugely changed(actually they got slightly better!), so it would be okay to use it.
One chip(A2-c1-b1) has very low inductance, so has large qubit frequency(7.66 GHz) and low coupling(<2 kHz). I don’t think it will work as a qubit.

The other chip(A2-g1-b1) is fine. The specs are

Qubit freq: 4.89 GHz
Qubit anharmonicity: 72.86 MHz

Readout freq: 6.211 GHz

Qubit-fundamental coupling : 85 kHz
Qubit-readout coupling : 3.6 MHz

Fundamental mode Q-factor: ~10^6
Readout mode Q-factor: ~10^5

#2 Updated by Daniil Frolov about 2 months ago

We found RI-2 cavity,
f= 2.985769709 GHz
Q ~ 1e8
temp of mixing chamber = 53 mK

#3 Updated by Daniil Frolov about 2 months ago

Some more accurate estimation of TI-2 Q from decay:
decay time from -50 dB to e times (green line) = 32 ms
Q = pi * 3E+9 * 32E-3 = 3E+8
Also attached raw data in "decay.csv"

#4 Updated by Daniil Frolov about 2 months ago

We found the resonance of RI-1,
freq = 2.985951150 GHz
Mixing chamber temperature = 65 mK (It is increased a little bit because we activated the switch.)

#5 Updated by Daniil Frolov about 2 months ago

estimation of TI-1 Q from decay (slightly better than TI-2)
decay time from -64 dB to e times (green line) = 39 ms
Q = pi * 3E+9 * 39E-3 = 3.67E+8
Attenuation through the cavity is ~15 dB higher.
Also attached raw data in "decay_RI-1.csv"

#6 Updated by Daniil Frolov about 2 months ago

CW measurement of chi_0 shift of fundamental mode of RI-2 cavity.
High power(20dBm, black)
Low power(-10dBm, blue)

chi_0 shift = 58.89 Hz


D:\redmine\runs\DR1_run_7\Labber\Data\2020\09\Data_0901\study_of_RI-2_TM010_3.hdf5

#7 Updated by Daniil Frolov about 2 months ago

Tried to test pulsed setup with new switches: measured Pappas' qubit readout. As expected, power levels went down by 10 dB because of new additional attenuation.

D:\redmine\runs\DR1_run_7\Labber\Data\2020\09\Data_0901\Pappas_PulsedReadout_trial.hdf5

#8 Updated by Roman Pilipenko 22 days ago

9/29/2020

DR warmed up and opened, as per e-mail from Alexander Romanenko from: Sep 26, 2020:

"...the problem is that the antenna coupling is too weak, and the resulting signal is too weak as well.

...6.19 GHz peak is floating around with minimal changes in temperature (even around 14-16 mK), that’s why it’s virtually impossible to measure its shift etc, as when you average and it moves around, it actually disappears.The same could be likely happening with the TM010 mode too...

...plan to change the antennas to stronger..."



Also available in: Atom PDF